How many moles of ions are in 25.5g of MgCl2
Answer #1
Molar mass of MgCl2 = 95.2 g/mol
Total number of moles of MgCl2 = mass / molar mass = 25.5 /95.2 = 0.268 mol
One mole of MgCl2 - 3 moles of ions Mg+ 2Cl-,
therefore no of moles of ions = 0.268 x 3 = 0.804 mol of ions
answered by: Language
Answer #2
molar mass of MgCl2 = 95.2 g/mol

moles of MgCl2 = mass / molar mass = 25.5 /95.2 = 0.268 mol
one mole of MgCl2 contains three moles of ions,
thus,
moles of ions = 0.268 x 3 = 0.804 mol

answered by: matin